# BS-LDM: Effective Bone Suppression in High-Resolution Chest X-Ray Images with Conditional Latent Diffusion Models

### v3

@@ 4 Conclusion
To enhance radiologists’ ability to detect lung lesions with greater precision and thoroughness, we introduce BS-LDM, an end-to-end framework for bone suppression in high-resolution CXR images, utilizing conditional LDMs. We utilize a multi-level hybrid loss-constrained VQGAN for perceptual compression to ensure detail preservation. To further refine the framework’s performance, we incorporate offset noise and a temporal adaptive thresholding strategy. These enhancements help reduce discrepancies in generating low-frequency information, thereby enhancing the quality and interpretability of the soft tissue images produced. Our framework effectively generates soft tissue images that achieve high levels of bone suppression while retaining essential details and preserving the integrity of lung lesions. Rigorous experimental and clinical evaluations confirm BS-LDM’s exceptional bone suppression capabilities and its significant clinical utility.

### v5

@@ 4 Conclusion
To enhance radiologists’ ability to detect lung lesions with greater precision and thoroughness, we introduce BS-LDM, an end-to-end framework for bone suppression in high-resolution CXR images, utilizing conditional LDMs. We utilize a multi-level hybrid loss-constrained VQGAN for perceptual compression to ensure detail preservation. To further refine the framework’s performance, we incorporate offset noise and a temporal adaptive thresholding strategy. These enhancements help reduce discrepancies in generating low-frequency information, thereby enhancing the quality and interpretability of the soft tissue images produced. Our framework effectively generates soft tissue images that achieve high levels of bone suppression while retaining essential details and preserving the integrity of lung lesions. Extensive experimental and downstream evaluations validate the superior bone suppression performance of BS-LDM and underscore its substantial clinical value. Although BS-LDM demonstrates potential, its clinical application is hindered by several obstacles. Key challenges include improving inference efficiency, optimizing memory usage, ensuring compatibility with the DICOM format, and refining interface design. Addressing these issues will be the focus of future efforts to enhance its viability for use in healthcare environments.